OpenStudy (anonymous):

What is the exact area of a circle with a radius of cm? A. 15.5 pi cm² B. 31 pi cm² C. 961 pi cm² D. 240.25 pi cm²

OpenStudy (ipwnbunnies):

\[Area of a circle = \pi r^2\] For an exact answer, keep pi as pi. :3

OpenStudy (anonymous):

ok so wat is it @iPwnBunnies

OpenStudy (ipwnbunnies):

Idek. Square the given radius, keep the pi the same. .-. Do you have a calculator?

OpenStudy (triciaal):

i can't read your value for r use the formula as above A= pi *r^2 pi is 22/7

OpenStudy (anonymous):

im so sorry its 31 @triciaal

OpenStudy (triciaal):

do not substitute for pi since pi is in the answer A = pi (31)^2 = 961 pi ans c